In a gumball machine there are 22 red, 14 orange, 9 green, and 15 white gumballs. You decide

to purchase a gumball from the machine.

a. What is the theoretical probability of getting an orange gumball? Express your answer as a

fraction in lowest terms


1
Expert's answer
2021-11-24T13:05:13-0500
"22+14+9+15=60"

a.

"P(orange)=\dfrac{14}{60}=\dfrac{7}{30}"